OpenStudy (anonymous):

find the equation of a line with a gradient of -2 that passes through the point 4 , -1

OpenStudy (amistre64):

theres nothing to find, they give you a slope, m, and a point (a,b). just use the point slope form of a line ... y = m(x-a)+b and simplify to your hearts content.

OpenStudy (anonymous):

use this equation ... y-y.=m(x-x.)................. eqtn of line.................. y-(-1)= -2(x-4)